i don't get how to do this at all:
it says to graph the equation:
6y+5 = 5y+3

Answer by mgmoeab(37)

The only thing you need to do is solve for 'y'
6y + 5 = 5y + 3
PUT LIKE-TERMS ON THE SAME SIDE OF THE EQUATION
6y - 5y = 3 - 5
y = -2 <--- you would only have to graph this line.
Horizontal line which passes through point (x, 2)